OpenStudy (anonymous):

8th grade and below! now answering questions! just type ur questions below!

OpenStudy (anonymous):

a+b+c=4 a^2+b^2+c^2=10 a^3+b^3+c^3=22 find the values of a, b ,and c

OpenStudy (anonymous):

ok one second also what grade is this?

OpenStudy (anonymous):

for the first sequence you can have 3 answers- 2+1+1, 1+3+0 and 2+2+0

OpenStudy (anonymous):

I googled one of the hardest equations to solve and thats what I sent you

OpenStudy (anonymous):

why would u do that...?

OpenStudy (anonymous):

i felt like messing with you.

OpenStudy (anonymous):

karmas a will get ya! have a great day! hope u learn something valuable at school instead of screwing with people!